OpenStudy (samigupta8):

The wavelengths involved in the spectrum of deuterium are slightly different from that of hydrogen spectrum because- 1).Size of two nuclei are different 2).Nuclear forces are different in two cases 3).Masses of two nuclei are different 4).Attraction between the electron and the nucleus is different ...

OpenStudy (samigupta8):

@priyar plss..help

OpenStudy (samigupta8):

@Michele_laino sir pls help

OpenStudy (samigupta8):

@Vincent-Lyon.Fr sir pls help

OpenStudy (mayankdevnani):

i think option C is correct !

OpenStudy (mayankdevnani):

because ENERGY DIFFERENCE is dependent on MASS OF NUCLEI.

OpenStudy (samigupta8):

Can u plss give the relation?

OpenStudy (samigupta8):

N energy difference between two energy states is given by Rz^2(1/n1^2-1/n2^2) in case of hydrogen like species spectrum....

OpenStudy (mayankdevnani):

R depends on MASS

OpenStudy (samigupta8):

Sure...???

OpenStudy (samigupta8):

Lemme check then....

OpenStudy (mayankdevnani):

sure

OpenStudy (samigupta8):

Value of R comes out to be 2mπ^2k^2e^4/h^2

OpenStudy (mayankdevnani):

i don't remember the exact value but i knew it depends on mass as you can see

OpenStudy (samigupta8):

Well , thanks @mayankdevnani it's correct the expression that i gave n the ans u proposed....

OpenStudy (mayankdevnani):

alright ! Thumbs Up :)

OpenStudy (michele_laino):

I think it is since the masses of the two nuclei are different. The problem in writing the energy leves of a system of a hydrogen like atom is a two-body problem, so it involves the consideration of the reduced \(\mu\) mass of such system: \[\huge \frac{1}{\mu } = \frac{1}{{{m_e}}} + \frac{1}{M}\] where \(m_e\) is the mass of the electron, and \(M\) is the mass of the nucleus of such hydrogen-like atom

OpenStudy (michele_laino):

of course, the reduced mass changes, when we change the hydrogen-like atom type